iridas or Component casts of capital) Compute the cost for the following sources of financing. company tax rate of 30% Debt that has a $100 par value (face value) and a contract or coupon interest rate of 11% A new issue would have issue costs of 5% of the $112.50 market value. The debt matures in 10 years. (b) A new ordinary share issue that paid an 18 cents dividend last year. Earnings per share have grown at a rate of 7% per year. This growth rate is expected to continue into the foresee- able future. The company maintains a constant dividend/earnings ratio of 30%. The price ofthis share is now $2.75, but 5% issue costs are anticipated. Dividends are fully franked for à taxation category 1 company) ie internal equity where the current market price of the ordinary shares is $4.30. The expected dividend this year should be 35 cents, increasing thereafter at a 7% annual growth rate Dividends are fully franked (for a taxation category 1 company). (d) Pref 13.5 of $1.75 erence shares for a taxation category 2 company paying an unfranked dividend of cents per share. If a new issue is offered, issue costs will be 12% of the current price
(e) A bond selling to yield 12% after issue costs, but prior to adjusting for the marginal com corpo- rate tax rate of 30%. In other words, 12% is the rate that equates the net proceeds bond with the present value of the future cash flows (principal and interest) fromthe
Answer #1

a. The pretax cost of debt is the yield to maturity which can be calculated as =RATE(nper,pmt,pv,fv) in excel where nper = 10, pmt =11%*100 =11, pv =112.50 and fv =100 + 5% issue cost =100+5 =105

Pre tax cost of debt =RATE(10,11,-112.50,105) =9.35%

After tax cost of debt = pretax cost * (1-tax rate)= 9.35*(1-0.30) = 6.55%

b. Cost of equity = D1/P0*(1-f) + g

D1 = 0.18*1.07 = 0.1926

P0 = 2.75

f = issue cost = 0.05

g = 7% = 0.07

Cost of equity = 0.1926/2.75*(1-0.05) +0.07

Cost of equity = 01437 = 14.37%

c. Cost of internal Equity = D1/P0 + g = 0.35/4.30 + 0.07 = 0.1514 = 15.14%

d. Cost of preference share = Dividend/ Price *(1- flatation cost) = 0.135/1.75*(1-0.12) = 0.08766

Cost of preference share = 8.77% (Rounded)